QS 22-27A Merchandising: Cash payments for merchandise LO P4

Torres Co. forecasts merchandise purchases of $14,800 in January, $17,600 in February, and $20,200 in March; 50% of purchases are paid in the month of purchase and 50% are paid in the following month. At December 31 of the prior year, the balance of Accounts Payable (for December purchases) is $28,000.

Prepare a schedule of cash payments for merchandise for each of the months of January, February, and March.QS 22-27A Merchandising: Cash payments for merchandise LO P4 Torres Co. forecasts merchandise purchases of $14,800 in January
Answer #1
January February March
Cash payment for current month (14,800*50%) = 7400 (17,600*50%) = 8,800 (20,200*50%) = 10,100
Cash payment for prior month 28,000 (14,800*50%) = 7400 (17600*50%) = 8800
Budgeted cash payments 35400 16200 18900

Comment if you face any issues